# Personalize to generalize: Towards a universal medical multi-modality generalization through personalization

### v1

@@ 5.2 Tuning for segmentation
Experimental settings. We utilize the AutoPET-II [15] dataset for segmentation, evaluating performance using the DICE metric. It is important to note that we employ the same training and testing splits as in Sec. 5.1 to avoid data leakage. Specifically, we adhere to the settings from the official challenge; DICE is calculated in the standard manner but is set to zero for false negatives and true negatives. Additionally, we introduce DICE- to include the mean across all samples, along with true positive rate (TPR), true negative rate (TNR), false negative rate (FNR), and false positive rate (FPR) for the missing modality segmentation evaluation. Our method is compared against nnUNET [21], UNETR [18], and SwinUNETR [17], which are trained directly on the dataset without pre-training. Notably, we also compare our approach with SwinUNETR using its pre-training strategy [40]. Please refer to training details in Appendix D.
Results. The full modality segmentation results are exhibited in Tab. 5. The results indicate that with proper model architecture, such as SwinUNETR, using both two modalities usually outperforms solely using PET. It can be observed that models using our pre-train improve the results across all metrics. Typically, SwinUNETR using our pre-train significantly exceeds it without our pre-trained model, indicating the personalized invariant learned by our pre-train generalizes to the downstream well and can boost the downstream tasks. Moreover, using our proposed components with the pre-train leads to the best DICE and DICE-. This validates that using the prior further emphasizes the personalized invariant, which yields the most segmentation improvements.
6 Special case: A complex scenario
@@
We introduce a more complex scenario, in which the pre-trained model for heterogeneous generalization settings is tuned downstream that span both heterogeneous and homogeneous generalization.
@@
Experimental settings. The pre-train model we adopted is from Sec. 5 that trained on AC-PET and CT. Specifically, we tune the model by using the Head and Neck Squamous Cell Carcinoma (HNCSS) dataset [16] as the training set and the Non-Small Cell Lung Cancer (NSCLC) dataset as the testing set. Both datasets are sourced from The Cancer Imaging Archive (TCIA) [9], and they contain paired non-attenuation-corrected PET (NAC-PET), attenuation-corrected PET (AC-PET), and CT scans. The model is pre-trained for heterogeneous generalization between AC-PET and CT. It is tuned for both homogeneous generalization between AC-PET and NAC-PET and heterogeneous generalization between NAC-PET and CT. Similar to the previous translation experiments, we use SSIM and PSNR as evaluation metrics. Performance in this scenario further validates the model’s generalization capabilities. Note here the training and testing data in the downstream task come from different domains. See training details in Appendix D.
@@ 7 Conclusion
This paper proposes a universal approach to address multi-modality generalization by approximating a personalized invariant representation, $\mathbb{X}_{h}$, through constraints of invariance, equivariance, and decomposition, guided by a learnable biological prior. We demonstrate that learning $\mathbb{X}_{h}$ is both feasible and highly beneficial for enhancing generalization in medical tasks. By discussing limitations, future directions, and social impact in Appendices B and C, our study may point to a promising path for achieving medical generalization through personalization in complex settings.
Appendix A Related work
@@
Appendix B Limitations, challenges, and future work
@@
Appendix C Social impact
@@
C.1 Downstream segmentation ablation study
@@
We demonstrate the effectiveness of our proposed components and discuss the process of learning an anatomy-invariant representation. Experimental results for downstream segmentation tasks and visualizations of the pre-trained models are presented in Tab. 7. All experiments are conducted under consistent settings to ensure a fair comparison.
Using all constraints together with $\mathbb{O}$ yields the best results. Consistent with Sec. 3.2.1, the results indicate that using different constraints alone may not guarantee improvements; however, incorporating all constraints along with $\mathbb{O}$ results in the best outcomes. This validates the plausibility of the $\mathbb{X}_{h}$ Hypothesis and demonstrates that achieving good approximation of it significantly enhances generalization.
Using prior $\mathbb{O}$ with decomposition constraint improves the model performance for different settings. Despite different settings, additionally using $\mathbb{O}$ with decomposition improves the downstream model performance. Combined with the improvements from modality transfer results in Tab. 4, it suggests that $\mathbb{O}$ helps with better obtaining anatomical structure.
The invariance and equivariance constraints can not be applied to the same feature. It needs to be highlighted that invariance and equivariance constraints can not be applied to the same features as they conflict with each other. As shown in task 3, without $\mathbb{O}$, invariance and equivariance constraints are applied to the latent feature simultaneously, leading to a significant performance drop. In comparison, apply equivariance constraint before using $\mathbb{O}$ and applying invariance constraint after using $\mathbb{O}$ yields the best results. This is because the geometrical transformation contained in $z^{i}_{h}$ needs to be accomplished by fetching other possible geometrical transformation information form $\mathbb{O}$ and then fused to be invariant.
Appendix D Experimental details
@@
D.1 Overall training procedure
We provide a pseudo-code for our approach. The loss calculation for Pre-training procedure is simplified as Algorithm 1 and Downstream tuning as Algorithm 2. It is notable that the empirical procedure is flexible as long as the $\mathbb{O}$ is properly used to construct $\mathbb{X}_{h}^{\prime}$ and those constraints are applied to $\mathbb{X}_{h}^{\prime}$.
@@ D.1 Overall training procedure
D.2 Homogeneous generalization: structural modalities in MRI
D.2.1 Pre-training and Modality transfer.
Experimental settings. We use four A100 GPUs for training. The learning rate we used for the modality transfer is set to $0.0002$, and the training epoch is set to $1000$. Both the number of input and out channels is set as $4$.
@@ D.2 Homogeneous generalization: structural modalities in MRI
D.2.2 Missing modality segmentation.
We use four A100 GPUs for tuning. The learning rate we used for the modality transfer is set to $0.0002$, and the training epoch is set to $1000$. Both the number of input and out channels is set as $4$.
@@
D.3 Heterogeneous generalization: PET and CT modalities
D.3.1 Modality transfer
@@ D.3.2 Downstream segmentation
D.4 Special case: Tuning from heterogeneous to homogeneous generalization with domain gap
@@
Appendix E More results
E.1 Modality transfer results on BRATS22:
Tab. 8 and Tab. 9 presents the generation result with standard derivations. The results of our method and SwinUNETR are produced by ourselves, while the rest of the results are gathered from [25]. Generated examples are presented in Figs. 6, 7 and 8.
@@ E.1 Modality transfer results on BRATS22:
E.2 Missing modality segmentation results on BRATS18:
@@
Figure 8: Generated images of our proposed method: slices across cerebellar hemisphere. Our method can generate defined cerebellar folia.

### v4

@@ 4.2.2 Fine-tuning stage: Segmentation
Experimental settings. We utilize the AutoPET-II [17] dataset for segmentation, evaluating performance using the DICE metric. It is important to note that we employ the same training and testing splits as in Sec. 4.2.1 to avoid data leakage. Specifically, we adhere to the settings from the official challenge; DICE is calculated in the standard manner but is set to zero for false negatives and true negatives. Additionally, we introduce DICE- to include the mean across all samples, along with true positive rate (TPR), true negative rate (TNR), false negative rate (FNR), and false positive rate (FPR) for the missing modality segmentation evaluation. Our method is compared against nnUNET [23], UNETR [20], and SwinUNETR [19], which are trained directly on the dataset without pre-training. Notably, we also compare our approach with SwinUNETR using its pre-training strategy [44]. Please refer to training details in Appendix F.
From scratch Dice
nnUnet [23] 33.1
SwinUNETR [19] 43.5
With 3D medical image SSL
SwinUNETR [44] 44.1
SwinMM [49] 44.2
PCRL v2 [59] 41.9
VoCo [51] 46.3
Ours (PUIR) 48.2
Table 6: Comparison between different SOTA medical 3D SSL methods on AUTOPET segmentation.
Results. The segmentation results are presented in Tab. 6, with additional details provided in Appendix Tab. 8. For a fair comparison, all other SSL methods were reproduced using the same pre-training and fine-tuning datasets as those used in our approach. Our approach significantly improves the DICE results in comparison to other SSL methods since we specifically address the cross-modality generalization. The performance gains on heterogeneous modalities also enhance the significance of our $\mathbb{X}_{h}$ hypothesis for heterogeneous generalization.
4.3 Fine-tuning special case: A complex scenario
@@
We introduce a more complex scenario in which the pre-trained model for heterogeneous generalization settings is tuned downstream to span both heterogeneous and homogeneous generalization.
@@
Experimental settings. The pre-train model we adopted is from Sec. 4.2 that trained on AC-PET and CT. Specifically, we tune the model by using the Head and Neck Squamous Cell Carcinoma (HNCSS) dataset [18] as the training set and the Non-Small Cell Lung Cancer (NSCLC) dataset as the testing set. Both datasets are sourced from The Cancer Imaging Archive (TCIA) [11], and they contain paired non-attenuation-corrected PET (NAC-PET), attenuation-corrected PET (AC-PET), and CT scans. The model is pre-trained for heterogeneous generalization between AC-PET and CT. It is tuned for both homogeneous generalization between AC-PET and NAC-PET and heterogeneous generalization between NAC-PET and CT. Similar to the previous translation experiments, we use SSIM and PSNR as evaluation metrics. Performance in this scenario further validates the model’s generalization capabilities. Note here the training and testing data in the downstream task come from different domains. See training details in Appendix F.
@@ 5 Conclusion
This paper proposes a universal approach to address multi-modality generalization by approximating a personalized invariant representation, $\mathbb{X}_{h}$, through constraints of invariance, equivariance, and decomposition, guided by a learnable biological prior. We demonstrate that learning $\mathbb{X}_{h}$ is both feasible and highly beneficial for enhancing generalization in medical tasks. Our method also correlates to personalized medicine, a transformative framework for $21^{st}$ century healthcare, tailoring medical treatments to each patient’s unique characteristics [50, 26, 5]. By discussing limitations, future directions, and social impact in Appendices C and D, our study may point to a promising path for achieving medical generalization through personalization in complex multi-modality medical analysis.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 92370119, No. 62376113, No. 62206225, No. 62436009, No. 62276258, No. 82394432, and No.92249302), the Shanghai Municipal Science and Technology Major Project (Grant No. 2023SHZDZX02), and XJTLU Funding REF-22-01-002. The computations in this research were performed using the CFFF platform of Fudan University.
Supplementary Material
Appendix A Brief descriptions of different modalities
Magnetic Resonance Imaging (MRI) scans [58]: use strong magnetic fields and radiofrequency currents yielding distinct sequences. Typically, MRI has different modalities, include T1, T2, T1ce and Flair.
Computed Tomography (CT) scans [36, 56] employ X-rays to measure its attenuation.
Positron Emission Tomography (PET) scans are expensive functional imaging scans that employ radiotracers emitting gamma rays to visualize and measure metabolic processes. Thus, PET scans have a large percentage of background areas.
Appendix B Related work
@@
Appendix C Limitations, challenges, and future work
@@
Appendix D Social impact
@@
Appendix E Downstream segmentation ablation study
@@
The effectiveness of our proposed components is demonstrated alongside an exploration of the methodology employed to develop an individual-invariant representation. Experimental results for downstream segmentation tasks and visualizations of the pre-trained models are presented in Tab. 7. All experiments are conducted under consistent settings to ensure a fair comparison.
Using all constraints together with $\mathbb{O}$ yields the best results. Consistent with Sec. 3.1.2, the results indicate that using different constraints alone may not guarantee improvements; however, incorporating all constraints along with $\mathbb{O}$ results in the best outcomes. This validates the plausibility of the $\mathbb{X}_{h}$ Hypothesis and demonstrates that achieving a good approximation of it significantly enhances generalization.
Using prior $\mathbb{O}$ with decomposition constraint improves the model performance for different settings. Despite different settings, additionally using $\mathbb{O}$ with decomposition improves the downstream model performance. Combined with the improvements from modality transfer results in Tab. 4, it suggests that $\mathbb{O}$ helps in better obtaining anatomical structure.
The invariance and equivariance constraints can not be applied to the same feature. It needs to be highlighted that invariance and equivariance constraints cannot be applied to the same features as they conflict with each other. As shown in task 3, without $\mathbb{O}$, invariance and equivariance constraints are applied to the latent feature simultaneously, leading to a significant performance drop. In comparison, applying equivariance constraint before using $\mathbb{O}$ and applying the invariance constraint after using $\mathbb{O}$ yields the best results. This is because the geometrical transformation contained in $z^{i}_{h}$ needs to be accomplished by fetching other possible geometrical transformation information from $\mathbb{O}$ and then fusing it to be invariant.
Appendix F Experimental details
@@
F.1 Overall training procedure
A pseudo-code is provided for our approach. The loss calculation for Pre-training procedure is simplified as Algorithm 1 and Downstream tuning as Algorithm 2. It is notable that the empirical procedure is flexible as long as the $\mathbb{O}$ is properly used to construct $\mathbb{X}_{h}^{\prime}$ and those constraints are applied to $\mathbb{X}_{h}^{\prime}$.
@@ F.1 Overall training procedure
F.2 Homogeneous generalization: structural modalities in MRI
F.2.1 Pre-training and Modality transfer.
Experimental settings. Four A100 GPUs are employed for training. The learning rate we used for the modality transfer is set to $0.0002$, and the training epoch is set to $1000$. Both the number of input and out channels is set as $4$.
@@ F.2 Homogeneous generalization: structural modalities in MRI
F.2.2 Missing modality segmentation.
Four A100 GPUs are employed for tuning. The learning rate we used for the modality transfer is set to $0.0002$, and the training epoch is set to $1000$. Both the number of input and out channels is set as $4$.
@@
F.3 Heterogeneous generalization: PET and CT modalities
F.3.1 Modality transfer
@@ F.3.2 Downstream segmentation
F.4 Fine-tuning special case: Tuning from heterogeneous to homogeneous generalization with domain gap
@@
Appendix G More results and analysis
G.1 More analysis on learnable biological prior
Figure 7: Visualizations of generated modalities with T1 as input of our method, which allows the capturing of subtle structures.
G.2 Segmentation results on AutoPET-II.
Detailed metrics results of AutoPET-II are presented in Tab. 8. The results indicate that with proper model architecture, such as SwinUNETR, using both two modalities usually outperforms solely using PET. It can be observed that models using our pre-train improve the results across all metrics. Typically, SwinUNETR using our pre-train significantly exceeds it without our pre-trained model, indicating the personalized invariant learned by our pre-train generalizes to the downstream well and can boost the downstream tasks. Moreover, using our proposed components with the pre-train leads to the best DICE and DICE-. This validates that using the prior further emphasizes the personalized invariant, which yields the most segmentation improvements.
Method Dice↑ Dice-↑ TPR↑ TNR↑ FNR↓ FPR↓
From scratch
nnUnet [23] 33.10 - - - - -
SwinUNETR [19] 43.45 62.60 90.32 62.73 9.68 37.27
SwinUNETR with different pre-train
With pre-train in [44] 44.06 57.79 89.25 73.64 10.75 26.36
With ours 48.20 61.16 88.17 77.27 11.83 22.72
G.3 Modality transfer results on BRATS22.
Tab. 9 and Tab. 10 presents the generation result with standard derivations. The results of our method and SwinUNETR are produced by ourselves, while the rest of the results are gathered from [27]. Generated examples are presented in Figs. 10, 11 and 12.
@@ G.1 More analysis on learnable biological prior
G.4 Missing modality segmentation results on BRATS18.
@@
G.5 Comparison with non-personalized methods.
We provide further visual evidence in Fig. 9 for an in-depth analysis. Fig. 9 shows that applying our constraints aligns CT and PET representations more closely, indicating smaller $d_{\mathcal{G}\Delta\mathcal{G}}(\mathcal{M}|\mathcal{H}^{\mathcal{S}},\mathcal{M}|\mathcal{H}^{\mathcal{U}})$ in Paper Eq. (10), yielding tightened bounds and better downstream performance (Dice:48.20), in comparison to the non-personalized baseline (Dice:43.5). This further supports our theoretical analysis.
Figure 9: TSNE of latent features of PET and CT images obtained under different applied constraints. Downstream performances are noted.
@@
Figure 12: Generated images of our proposed method: slices across the cerebellar hemisphere. Our method can generate defined cerebellar folia.